one-fourth of eric age last year plus triple his age next year is 45. How old is Eric?
Answer by mananth(16946)

Eric's age now------x
last year x-1 years

next year age = x+1
(x-1)/4 + 3(x+1) =45
x-1 +12(x+1)=4*45
x-1+12x+12 =180
13x+11=180
13x= 180-11
13x=169
x= 13 years age now